OpenStudy (anonymous):

how to find argument of COMPLEX NUMBERS? take this question-Convert the given complex number in polar form: 1 – i

ganeshie8 (ganeshie8):

\[\large a + ib\] \[\large \text{argument} = \tan^{-1}\left(\frac{b}{a}\right)\]

OpenStudy (anonymous):

 now tell me this method sir